# Convergence of Schrödinger operators on domains with scaled resonant potentials

### v1

@@ 4.3. No uniform ellipticity
We finally show that if Lemma 4.6 is not true then $(\mathfrak{a}_{\varepsilon})_{\varepsilon}$ is not uniformly elliptic:
@@
[Proof]Without loss of generality we can assume that $\omega>0$ in the definition of uniform ellipticity. If $(\mathfrak{a}_{\varepsilon})_{\varepsilon}$ was uniformly elliptic, then for any $\hat{v}\in\ring{H}^{1}(\Omega)$
@@ 5. Absence of norm resolvent convergence
Let $V\in C^{\infty}_{\mathrm{c}}(\overline{{{\mathbb{R}}}}_{+})$ such that $\operatorname{supp}V\subset[0,a]$ with $a>0$ be a non-positive resonant potential in the sense of Definition 1.1; i.e. the self-adjoint Schrödinger operator $L_{2}((0,a))$ corresponding to the quadratic form $f\mapsto\int_{0}^{a}(\lvert f^{\prime}\rvert^{2}+V\lvert f\rvert^{2})\operatorname{d\!}x$ with domain $\{f\in H^{1}((0,a))\colon f(0)=0\}$ has eigenvalue zero. Recall that there exists a sequence of real numbers $\{\alpha_{n}\}_{n\in{\mathbb{N}}}$, $1=\alpha_{1}<\alpha_{2}<\alpha_{3}<\dots<\alpha_{n}<\dots$ such that $\alpha_{n}\to\infty$, for which the multiple $\alpha_{n}V$ of the potential $V$ is resonant for all $n\in{\mathbb{N}}$. It remains to note that for all $n\in{\mathbb{N}}$ sufficiently large the resonant potential $\alpha_{n}V$ necessarily satisfies Hypothesis 5.1. Thus, the family of resonant potentials satisfying the above hypothesis is non-void.
@@
$\mathfrak{a}_{\varepsilon}^{(m)}[f]$ :$\vcentcolon=\int_{0}^{1}\Bigl{(}\lvert f^{\prime}(r)\rvert^{2}+\frac{m^{2}}{r^{2}}\lvert f(r)\rvert^{2}+\frac{1}{\varepsilon^{2}}V\bigg{(}\frac{1-r}{\varepsilon}\bigg{)}\lvert f(r)\rvert^{2}\Bigr{)}r\operatorname{d\!}r,$
$\operatorname{dom}\mathfrak{a}_{\varepsilon}^{(m)}$ :=$\vcentcolon=\big{\{}f\in L_{2}((0,1);r\operatorname{d\!}r)\colon\mathfrak{a}_{\varepsilon}^{(m)}[f]<\infty\big{\}}.$
The symmetric quadratic form $\mathfrak{a}_{\varepsilon}^{(m)}$ is closed, densely defined, and lower-semibounded for any $m\in{\mathbb{Z}}$. The mentioned properties of $\mathfrak{a}_{\varepsilon}^{(m)}$ follow immediately from the perturbation result [K, Chapter VI, Theorem 1.33] and the fact that this form can be represented as a sum of a bounded quadratic form
@@
[Proof]Recall that the Robin Laplacian ${\mathsf{A}}$ is bounded from below. Let us choose $\beta<0$ such that $\beta<\inf\sigma({\mathsf{A}})$. By Lemma 5.3 we can find $m_{1}\in{\mathbb{N}}$ and $\varepsilon_{1}>0$ such that $\lambda_{1}^{(m_{1})}(\varepsilon_{1})=\beta$. By item (c) of the same lemma we can choose integer $m_{2}>m_{1}$ such that $\lambda_{1}^{(m_{2})}(\varepsilon_{1})\geq 0$. Hence, by items (a) and (b) of Lemma 5.3 we can find $\varepsilon_{2}\in(0,\varepsilon_{1})$ such that $\lambda_{1}^{(m_{2})}(\varepsilon_{2})=\beta$. Analogously, we can find $m_{3}>m_{2}$ and $\varepsilon_{3}\in(0,\varepsilon_{2})$ such that $\lambda_{1}^{(m_{3})}(\varepsilon_{3})=\beta$. Thus, repeating the construction, we conclude that there exists sequences of real numbers $\varepsilon_{1}>\varepsilon_{2}>\dots>\varepsilon_{k}>\dots>0$ and integers $m_{1}<m_{2}<\dots<m_{k}<\dots<+\infty$ such that $\lambda_{1}^{(m_{k})}(\varepsilon_{k})=\beta$ for all $k\in{\mathbb{N}}$. Moreover, it follows from Lemma 5.3 (c) that $\varepsilon_{k}\leq\frac{1}{m_{k}}\sqrt{\lVert V\rVert_{\infty}}\to 0$ as $k\to\infty$
@@ Acknowledgements
The first named author (VL) is grateful to Casa Mathematica Oaxaca (CMO) for providing the possibility to participate in the workshop “Analytic and Geometric Aspects of Spectral Theory” in August 2022, where this project was initiated and to University of Trier for funding a research stay in February 2023, during which a part of the results presented in this manuscript were obtained. The research was supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 873071.
@@
where $f_{\mu}\in\ring{H}^{1}({\mathbb{R}}_{+})\cap H^{2}({\mathbb{R}}_{+})$ satisfies the differential equation $-f_{\mu}^{\prime\prime}+Vf_{\mu}=\mu f_{\mu}$ and where $\mu<0$ is as in Hypothesis 5.1. Since $\operatorname{supp}V\subset[0,a]$, we obtain that for $t>a$
@@
(c) The statement is a consequence of the representation of the quadratic form $\mathfrak{a}_{\varepsilon}^{(m)}$ and the fact that under the assumption $\varepsilon\geq(1/\lvert m\rvert)\sqrt{\lVert V\rVert_{\infty}}$ the function on $(0,1)$ acting as

### v2

@@ 4.3. No uniform ellipticity
We finally show that if the estimate in Lemma 4.6 with any constant is not true then $(\mathfrak{a}_{\varepsilon})_{\varepsilon}$ is not uniformly elliptic:
@@
[Proof]Without loss of generality we can assume that $\omega>0$ in the definition of uniform ellipticity. If $(\mathfrak{a}_{\varepsilon})_{\varepsilon}$ was uniformly elliptic, then for any $\hat{v}\in\operatorname{dom}\mathsf{A}_{\varepsilon}$
@@ 5. Absence of norm resolvent convergence
Let $V\in C^{\infty}_{\mathrm{c}}(\overline{{{\mathbb{R}}}}_{+})$ such that $\operatorname{supp}V\subset[0,a]$ with $a>0$ be a non-positive resonant potential in the sense of Definition 1.1; i.e. the self-adjoint Schrödinger operator in $L_{2}((0,a))$ corresponding to the quadratic form $f\mapsto\int_{0}^{a}(\lvert f^{\prime}\rvert^{2}+V\lvert f\rvert^{2})\operatorname{d\!}x$ with domain $\{f\in H^{1}((0,a))\colon f(0)=0\}$ has eigenvalue zero. Recall that there exists a sequence of real numbers $\{\alpha_{n}\}_{n\in{\mathbb{N}}}$, $1=\alpha_{1}<\alpha_{2}<\alpha_{3}<\dots<\alpha_{n}<\dots$ such that $\alpha_{n}\to\infty$, for which the multiple $\alpha_{n}V$ of the potential $V$ is resonant for all $n\in{\mathbb{N}}$. It remains to note that for all $n\in{\mathbb{N}}$ sufficiently large the resonant potential $\alpha_{n}V$ necessarily satisfies Hypothesis 5.1. Thus, the family of resonant potentials satisfying the above hypothesis is non-void.
@@
$\mathfrak{a}_{\varepsilon}^{(m)}[f]$ $\vcentcolon=\int_{0}^{1}\Bigl(\lvert f^{\prime}(r)\rvert^{2}+\frac{m^{2}}{r^{2}}\lvert f(r)\rvert^{2}+\frac{1}{\varepsilon^{2}}V\bigg(\frac{1-r}{\varepsilon}\bigg)\lvert f(r)\rvert^{2}\Bigr)r\operatorname{d\!}r,$
$\operatorname{dom}\mathfrak{a}_{\varepsilon}^{(m)}$ $\vcentcolon=\big\{f\in L_{2}((0,1);r\operatorname{d\!}r)\colon\mathfrak{a}_{\varepsilon}^{(m)}[f]<\infty\big\}.$
The symmetric quadratic form $\mathfrak{a}_{\varepsilon}^{(m)}$ is closed, densely defined, and lower-semibounded in $L^{2}((0,1);r\operatorname{d\!}r)$ for any $m\in{\mathbb{Z}}$. The mentioned properties of $\mathfrak{a}_{\varepsilon}^{(m)}$ follow immediately from the perturbation result [K, Chapter VI, Theorem 1.33] and the fact that this form can be represented as a sum of a bounded quadratic form
@@
[Proof]Recall that the Robin Laplacian $\mathsf{A}$} is bounded from below. Let us choose $\beta<0$ such that $\beta<\inf\sigma({\mathsf{A}})$. By Lemma 5.3 we can find $m_{1}\in{\mathbb{N}}$ and $\varepsilon_{1}>0$ such that $\lambda_{1}^{(m_{1})}(\varepsilon_{1})=\beta$. By item (c) of the same lemma we can choose integer $m_{2}>m_{1}$ such that $\lambda_{1}^{(m_{2})}(\varepsilon_{1})\geq 0$. Hence, by items (a) and (b) of Lemma 5.3 we can find $\varepsilon_{2}\in(0,\varepsilon_{1})$ such that $\lambda_{1}^{(m_{2})}(\varepsilon_{2})=\beta$. Analogously, we can find $m_{3}>m_{2}$ and $\varepsilon_{3}\in(0,\varepsilon_{2})$ such that $\lambda_{1}^{(m_{3})}(\varepsilon_{3})=\beta$. Thus, repeating the construction, we conclude that there exists sequences of real numbers $\varepsilon_{1}>\varepsilon_{2}>\dots>\varepsilon_{k}>\dots>0$ and integers $m_{1}<m_{2}<\dots<m_{k}<\dots<+\infty$ such that $\lambda_{1}^{(m_{k})}(\varepsilon_{k})=\beta$ for all $k\in{\mathbb{N}}$. Moreover, it follows from Lemma 5.3 (c) that $\varepsilon_{k}\leq\frac{1}{m_{k}}\sqrt{\lVert V\rVert_{\infty}}\to 0$ as $k\to\infty$.
@@ Acknowledgements
The first named author (VL) is grateful to Casa Mathematica Oaxaca (CMO) for providing the possibility to participate in the workshop “Analytic and Geometric Aspects of Spectral Theory” in August 2022, where this project was initiated and to University of Trier for funding a research stay in February 2023, during which a part of the results presented in this manuscript were obtained. The research was supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 873071. We are grateful to the anonymous referees for valuable suggestions.
@@
where $f_{\mu}\in\ring{H}^{1}({\mathbb{R}}_{+})\cap H^{2}({\mathbb{R}}_{+})$ satisfies the differential equation $f_{\mu}^{\prime\prime}+Vf_{\mu}=\mu f_{\mu}$ and where $\mu<0$ is as in Hypothesis 5.1. Since $\operatorname{supp}V\subset[0,a]$, we obtain that for $t>a$
@@
(c) The statement is a consequence of the representation of the quadratic form $\mathfrak{a}_{\varepsilon}^{(m)}$ and the fact that under the assumption $\varepsilon\geq(1/\lvert m\rvert)\sqrt{\lVert V\rVert_{\infty}}$ the function on $(0,1)$ given by